OpenStudy (anonymous):

Karen earns $65,000 a year as an optometrist. She received a bonus of b dollars last year and expects to get double that amount as a bonus this year. Write an algebraic expression to show the total amount Karen expects to earn this year.

OpenStudy (anonymous):

6500+b=2b

OpenStudy (anonymous):

wait....

OpenStudy (anonymous):

its 65000+b=65000+2b

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

no prob!